# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ B.1 Moment Bounds
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
@@
Using the definition of $\mathsf{P}$ in (1.7), it holds that
@@
where application of Fubini’s theorem yields the last equality. The first inequality in (B.1) then follows from Assumption (H2). For the second inequality in (B.1), we first note that for any $a,v\in^{d_{u}}$ it holds that
$\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u$
$\qquad\succcurlyeq\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u;$ (B.2)
this follows from Lemma A.1 by noting that $\Psi(v)$ is the mean under the Gaussian probability measure $\mathcal{N}\bigl{(}\Psi(v),\Sigma\bigr{)}.$ Letting now $a$ be the mean under the measure $\mathsf{P}\mu$, we observe that, by conditioning on $v$ and using (B.1),
$\mathcal{C}(\mathsf{P}\mu)$ $=\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\,\mathsf{P}\mu(u)\,\mathrm{d}u$
$=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}(u-a)\otimes(u-a)\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$\succcurlyeq\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\left(\int_{{}^{d_{u}}}\bigl{(}u-\Psi(v)\bigr{)}\otimes\bigl{(}u-\Psi(v)\bigr{)}\exp\Bigl{(}-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\Bigr{)}\,\mathrm{d}u\right)\,\mu(\mathrm{d}v)$
$=\int_{{}^{d_{u}}}\Sigma\,\mu(\mathrm{d}v)=\Sigma,$
hence $\mathcal{C}(\mathsf{P}\mu)\succcurlyeq\Sigma$. On the other hand, using Lemma A.1 again and noting that, by the Cauchy-Schwarz inequality, $ww^{\mathsf{T}}\preccurlyeq(w^{\mathsf{T}}w)I_{d_{u}}$ for any vector $w\in^{d_{u}}$, we deduce that
@@
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumption H, it holds that
@@
The inequalities (B.4a) and (B.4b) follow from Assumption H and the fact that
@@
We now turn our attention to establishing inequality (B.5a). To this end, let $\phi\colon^{d_{u}}\to^{d_{u}+d_{y}}$ define the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, and let $\phi_{\sharp}$ denote the associated pushforward map on measures. We note that
@@
It holds for any $(a,b)\in^{d_{u}}\times^{d_{y}}$ that
$2\begin{pmatrix}aa^{\mathsf{T}}&0_{d_{u}\times d_{y}}\\ 0_{d_{y}\times d_{u}}&bb^{\mathsf{T}}\end{pmatrix}-\begin{pmatrix}a\\ b\end{pmatrix}\otimes\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}a\\ -b\end{pmatrix}\otimes\begin{pmatrix}a\\ -b\end{pmatrix}\succcurlyeq 0_{(d_{u}+d_{y})\times(d_{u}+d_{y})}.$
Hence, it follows that
@@
which enables to deduce, using Lemmas A.1 and B.1 and Assumption H, that
@@
where we used (B.3) and the calculation resulting in (B.7a) in the last inequality. Noting this inequality in combination with (B.8) yields the upper bound (B.5a). To establish the lower bound (B.5b), we first note that by the Cauchy–Schwarz inequality, for any $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
@@
$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b}.$
Hence, by Young’s inequality, we deduce that for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
@@
where we applied (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}I_{d_{y}}$ in the last inequality. We then use (B.8) to deduce that
@@
Taking $\varepsilon=\frac{2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}$, so that the coefficient of $\lvert b\rvert^{2}$ is $\frac{\gamma}{2}$, we obtain
@@
In order to finally obtain (B.5b), we use (B.7b) to deduce that
@@
which concludes the proof. ∎
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (1.11) initialized at the probability measure $\mu_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th order polynomial moment $\mathcal{M}_{q}(\mu_{0})<\infty$. If Assumption H holds then there exists $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$ such that
@@
where (B.13a) follows from Fubini’s theorem, the inequality (B.13b) from properties of Gaussians, and (B.13c) from application of Assumption H. We note that in (B.13c) the constant $C$ depends on $\kappa_{\Psi},\Sigma$. Now, to obtain a lower bound on the denominator of (B.12), we observe that
@@
Assume that the probability measures $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (2.4) initialized at the probability measure $\mu^{\rm EK}_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th polynomial order moment $\mathcal{M}_{q}\bigl{(}\mu_{0}^{\rm EK}\bigr{)}<\infty$. If Assumption H holds then there exists $C=C\Bigl{(}\mathcal{M}_{q}\bigl{(}\mu_{0}^{\rm EK}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ such that
@@
For $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ with finite second moments, it holds that
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma 20], where a complete proof is given. ∎
@@ B.2 Action of $\mathsf{T}_{j}$ on Gaussians
Fix $y^{\dagger}_{j+1}\in^{d_{y}}$ and let $\pi$ be a Gaussian measure over ×${}^{d_{u}}\times^{d_{y}}$. Then the probability measure $\mathsf{B}_{j}\pi$, with $\mathsf{B}_{j}$ defined in (1.10), coincides with the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$, where $\mathscr{T}$ is defined in (2.3).
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma 21], where a complete proof is given. ∎
@@ B.3 Stability Results
Suppose that $\Psi$ and $\Sigma$ satisfy Assumptions (H2) and (H5), respectively. Then it holds that
@@
which concludes the proof. ∎
@@
Suppose that $h$ and $\Gamma$ satisfy Assumptions (H3) and (H5), respectively. Then it holds that
@@
Under Assumption H, for any probability measure $\mu\in\mathcal{P}(^{d_{u}})$ with $\mathcal{M}_{2}(\mu)<\infty$ there exists a constant $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}>0$ such that
@@
For ease of exposition we write $y^{\dagger}=y^{\dagger}_{j+1}$. We define the $y$-marginal densities
@@
It holds that
@@
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)=\Gamma+\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\otimes h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}-\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]}\otimes\mathsf{P}\mu\bigl{[}h(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})\bigr{]},$ (B.19)
which by Assumption H is bounded from below by $\Gamma$ and from above by $\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}$. To deduce this we have again made use of the fact that $aa^{\mathsf{T}}\preccurlyeq(a^{\mathsf{T}}a)I_{d}$ for any vector $a\in^{d_{u}}$, by the Cauchy-Schwarz inequality. Therefore, noting that (B.4b) implies that $\left\lvert\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\right\rvert\leqslant C$ for a constant $C$ depending only on the parameters $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma$, we obtain
@@
The function $\beta_{\mu}$ can be bounded from below using similar reasoning. Indeed, by Assumption H we have that for all $y\in^{d_{y}}$,
@@
$\geqslant\frac{C\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}-\left\lVert\Gamma^{-1}\right\rVert\kappa^{2}_{h}\int_{{}^{d_{u}}}|u|^{2}\mathsf{P}\mu\mathrm{d}u\Bigr{)}}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}$ (B.21a)
@@
where we applied Jensen’s inequality in (B.21a) and the constant in (B.21b) depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.
@@
We begin by noting that for fixed $u\in^{d_{u}}$, the functions $y\mapsto\mathsf{Q}\mathsf{P}\mu(u,y)$ and $y\mapsto\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)$ are Gaussians up to constant factors. The covariance matrix of the former is $\Gamma$; it is possible to deduce that the covariance of the latter is given by
@@
where we have used the formula for the covariance of the conditional distribution of a Gaussian. We note that $\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ is positive semi-definite, hence by (B.19) and its upper bound it follows that the matrix (B.22) is bounded from above by ${2}\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}+\Gamma$. With the same notation as in (B.8), we find that
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)-\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ (B.23)
$\qquad=\Gamma+\Bigl{(}\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)-\mathcal{C}^{yu}(\phi_{\sharp}\mathsf{P}\mu)\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\phi_{\sharp}\mathsf{P}\mu)\Bigr{)}\succcurlyeq\Gamma;$ (B.24)
here, to deduce that the term in brackets is positive semi-definite, we observe that the term is the Schur complement of the block $\mathcal{C}^{uu}(\phi_{\sharp}\mathsf{P}\mu)$ of the matrix $\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$. Hence, (B.22) is bounded from below by $\Gamma$. Therefore the integral in the first term of (B.18) may be bounded from above by applying the result of Lemma A.2 in Appendix A with parameter $\alpha=\alpha\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma\bigr{)}$, yielding
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u$
$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad\leqslant C\int_{{}^{d_{y}}}\int_{{}^{d_{u}}}\left(1+|u|^{2}+|y|^{2}\right)\bigl{\lvert}\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\bigr{\rvert}\,\mathrm{d}u\,\mathrm{d}y$
$\qquad=Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu),$ (B.25)
where $C$ depends on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$.

### v2

@@ B.1 Moment Bounds
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumptions (H3) and (H2), it holds that
@@
The proof of this lemma follows the steps of the proof of [carrillo2022ensemble, Lemma B.1]; however, here different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper are required. Using the definition of $\mathsf{P}$ in (1.7a), it holds that
@@
where application of Fubini’s theorem yields the last equality. The first inequality in (B.1) then follows from Assumption (H3). Obtaining the lower bound of the second inequality in (B.1) may be done identically to the proof of [carrillo2022ensemble, Lemma B.1]. We turn our attention to obtaining the upper bound. Using Lemma A.1 and noting that $ww^{\mathsf{T}}\preccurlyeq(w^{\mathsf{T}}w)I_{d_{u}}$ for any vector $w\in^{d_{u}}$ by the Cauchy-Schwarz inequality, we deduce that
@@
Let $\mu$ be a probability measure on ${}^{d_{u}}$ with bounded first and second order polynomial moments $\mathcal{M}_{1}(\mu),\mathcal{M}_{2}(\mu)<\infty$. Under Assumptions (H3), (H4) and (H2), it holds that
@@
The inequalities (B.3a) and (B.3b) follow from the assumptions and the fact that
@@
To obtain the covariance bounds, we proceed using analogous steps to the proof of [carrillo2022ensemble, Lemma B.2]; however, here different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper are required. We begin by establishing inequality (B.4a). To this end, letting $\phi\colon^{d_{u}}\to^{d_{u}+d_{y}}$ define the map $\phi(u)=\bigl{(}u,h(u)\bigr{)}$, it holds that
@@
As shown in the proof of [carrillo2022ensemble, Lemma B.2] we have, for all $(a,b)\in^{d_{u}}\times^{d_{y}}$,
@@
from which we establish, using Lemmas A.1 and B.1 and Assumptions (H3), (H4) and (H2), that
@@
where we applied (B.2) and the calculation resulting in (B.6a) in the last inequality. Noting this inequality in combination with (B.7) yields the upper bound (B.4a). We now show that
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}\geqslant(1-\varepsilon)\sigma\left\lvert a\right\rvert^{2}+\left(\gamma-\left(\frac{1}{\varepsilon}-1\right)\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}\right)\lvert b\rvert^{2}.$ (B.9)
in order to establish the lower bound (B.4b). The argument is similar to that in [carrillo2022ensemble, Lemma B.2] but is explicit in certain constants whose dependencies on moments need to be controlled in this paper. We first note that for any $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and all $(a,b)\in^{d_{u}}\times^{d_{y}}$ it holds that
@@
$\leqslant\sqrt{a^{\mathsf{T}}\mathcal{C}^{uu}(\pi)a}\,\sqrt{b^{\mathsf{T}}\mathcal{C}^{yy}(\pi)b},$
by the Cauchy–Schwarz inequality. Hence, we deduce that for all $\varepsilon\in(0,1)$ and for all $(a,b)\in^{d_{u}}\times^{d_{y}}$
@@
where we applied Young’s inequality for the bound in the first line, and (B.1) and the bound $\mathcal{C}^{yy}(\phi_{\sharp}\mathsf{P}\mu)\preccurlyeq\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}I_{d_{y}}$ for the bound in the second line. We then apply (B.7) to find
@@
Choosing $\varepsilon=\frac{2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}{\gamma+2\mathsf{P}\mu\bigl{[}|h|^{2}\bigr{]}}$, so that the coefficient of the $\lvert b\rvert^{2}$ term is $\frac{\gamma}{2}$, we obtain
@@
In order to deduce (B.4b), we use (B.6b) to find that
@@
from which we conclude the desired result. ∎
@@
Let $q\geqslant 2$ be an integer. Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (1.9) initialized at the probability measure $\mu_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th order polynomial moment $\mathcal{M}_{q}(\mu_{0})<\infty$. If Assumptions (H1), (H3), (H4) and (H2) hold then there exists $C=C\bigl{(}\mathcal{M}_{q}(\mu_{0}),J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$ such that
@@
where (B.13a) follows from Fubini’s theorem, the inequality (B.13b) from properties of Gaussians and Assumption (H2), and (B.13c) from application of Assumption (H3). We note that in (B.13c) the constant $C$ depends on $\kappa_{\Psi},\Sigma$. Now, to obtain a lower bound on the denominator of (B.12), we observe that
@@
Let $q\geqslant 2$ be an integer. Assume that the probability measures $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained from the dynamical system (2.4) initialized at the probability measure $\mu^{{\rm EK}}_{0}\in~{}\mathcal{P}(^{d_{u}})$ with bounded $q$th polynomial order moment $\mathcal{M}_{q}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)}<\infty$. If Assumptions (H1), (H3), (H4) and (H2) hold then there exists $C=C\Bigl{(}\mathcal{M}_{q}\bigl{(}\mu_{0}^{{\rm EK}}\bigr{)},J,\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\Bigr{)}$ such that
@@
Let $\mu_{1},\mu_{2}\in\mathcal{P}(^{n})$ have finite second moments, then the following bounds hold:
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma B.4], where a complete proof is given. ∎
@@ B.2 Action of $\mathsf{T}_{j}$ on Gaussians
Fix $y^{\dagger}_{j+1}\in^{d_{y}}$ and let $\pi$ be a Gaussian measure over ×${}^{d_{u}}\times^{d_{y}}$. Then the probability measure $\mathsf{B}_{j}\pi$, with $\mathsf{B}_{j}$ defined in (1.8b) is equivalent to the probability measure $\mathsf{T}_{j}\pi=\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,y^{\dagger}_{j+1})_{\sharp}\pi$, where $\mathscr{T}$ is defined in (2.3b).
@@
The lemma as stated may be found in [carrillo2022ensemble, Lemma B.5], where a complete proof is given. ∎
@@ B.3 Stability Results
Suppose that $\Sigma$ and $\Psi$ satisfy Assumptions (H3) and (H2), respectively. Then it holds that
@@
thus concluding the proof. ∎
@@
Suppose that $\Gamma$ and $h$ satisfy Assumptions (H4) and (H2), respectively. Then it holds that
@@
Suppose that Assumptions (H1), (H3), (H4) and (H2) are satisfied. Then for any probability measure $\mu\in\mathcal{P}(^{d_{u}})$ with $\mathcal{M}_{2}(\mu)<\infty$ there exists a constant $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}>0$ such that
@@
The proof of this lemma follows the steps of the proof of [carrillo2022ensemble, Lemma B.9]; however, each step requires different bounds reflecting the assumptions on $\Psi$ and $h$ in this paper. For ease of exposition we write $y^{\dagger}=y^{\dagger}_{j+1}$. We define the $y$-marginal densities
@@
By definition of $\mathsf{B}_{j}$, it holds that
@@
$\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)=\Gamma+\mathsf{P}\mu\bigl{[}h\otimes h\bigr{]}-\mathsf{P}\mu\bigl{[}h\bigr{]}\otimes\mathsf{P}\mu\bigl{[}h\bigr{]}.$ (B.19)
Clearly $\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)\succcurlyeq\Gamma$. Furthermore, it holds that $\mathcal{C}^{yy}(\mathsf{Q}\mathsf{P}\mu)\preccurlyeq\Gamma+2\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}$ by (B.8). Therefore, noting that (B.3b) implies that $\left\lvert\mathcal{M}^{y}(\mathsf{Q}\mathsf{P}\mu)\right\rvert\leqslant C$ for a constant $C$ depending only on the parameters $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma$, we obtain
@@
The function $\beta_{\mu}$ can be bounded from below using similar reasoning. Indeed, by Assumptions (H3), (H4) and (H2) we have that for all $y\in^{d_{y}}$,
@@
$\geqslant\frac{1}{\sqrt{(2\pi)^{d_{y}}\det(\Gamma)}}\exp\Bigl{(}-\left\lVert\Gamma^{-1}\right\rVert|y|^{2}-\left\lVert\Gamma^{-1}\right\rVert\kappa^{2}_{h}\int_{{}^{d_{u}}}\bigl{(}2+|u|^{2}\bigr{)}\mathsf{P}\mu(u)\,\mathrm{d}u\Bigr{)}$ (B.21a)
@@
where we applied Jensen’s inequality in (B.21a) and the constant in (B.21b) depends on $\mathcal{M}_{2}(\mu),\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
@@
First note that, for fixed $u\in^{d_{u}}$, the functions $\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ and $\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ are Gaussians up to constant factors, with covariance matrices given respectively by $\Gamma$ and
@@
where we have used the formula for the covariance of the conditional distribution of a Gaussian. We note that $\mathcal{C}^{yu}(\mathsf{Q}\mathsf{P}\mu)\mathcal{C}^{uu}(\mathsf{Q}\mathsf{P}\mu)^{-1}\mathcal{C}^{uy}(\mathsf{Q}\mathsf{P}\mu)$ is positive semi-definite, hence by (B.19) and its upper bound it follows that the matrix (B.22) is bounded from above by ${2}\kappa_{h}^{2}\bigl{(}1+\operatorname*{tr}(\Sigma)+2\kappa_{\Psi}^{2}\bigl{(}1+\mathcal{M}_{2}(\mu)\bigr{)}\bigr{)}I_{d_{y}}+\Gamma$. As shown in the proof of [carrillo2022ensemble, Lemma B.9], the matrix (B.22) is bounded from below by $\Gamma$ (see [carrillo2022ensemble, Equation (B.20)]). It follows from using Lemma A.2 in Appendix A, with parameter $\tau=\tau\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma\bigr{)}$, that
$\int_{{}^{d_{u}}}\left(1+|u|^{2}\right)\left|\mathsf{G}\mathsf{Q}\mathsf{P}\mu(u,y)-\mathsf{Q}\mathsf{P}\mu(u,y)\right|\mathrm{d}u\leqslant Cd_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}\mu,\mathsf{Q}\mathsf{P}\mu),$ (B.23)
where $C$ is a constant depending on $\mathcal{M}_{2}(\mu),\kappa_{h},\kappa_{\Psi},\Sigma,\Gamma$. We refer to [carrillo2022ensemble, Equation (B.21)] for the detailed steps used to establish this bound.